OpenStudy (anonymous):

The area of two sphere have the ratio 4:25. what is the ratio of the radii?

OpenStudy (amistre64):

how do you determine the area of a sphere?

OpenStudy (anonymous):

Sorry, but I don't know how ..

pooja195 (pooja195):

\[\huge~\rm~A=\pi~r^2\]

OpenStudy (anonymous):

oh wait can i use the formula for sa surface area of the sphere?

pooja195 (pooja195):

@amistre64

OpenStudy (anonymous):

4(4piR^2)=25(4piR^2) like this?

OpenStudy (anonymous):

area of surface of sphere=4 pi r^2

OpenStudy (anonymous):

\[\frac{ 4 \pi r^2 }{ 4 \pi R^2 }=\frac{ 4 }{ 25 }\] \[\frac{ r }{ R }=?\]

OpenStudy (anonymous):

and then?

OpenStudy (amistre64):

if r^2 = 4 and R^2 = 25 what are the values of r and R ?

OpenStudy (amistre64):

as far as ratios go, there is a property: quadratics are the square of linears radius is a linear measure, area is a quadratic measure.

OpenStudy (anonymous):

\[\frac{ r^2 }{ R^2 }=\frac{ 4 }{ 25 },\frac{ r }{ R }=\sqrt{\frac{ 4 }{ 25 }}=?\]

OpenStudy (anonymous):

and the ratio is 2:5...

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I did not realize that I can cancel the 4 pi and squared both sides haha